# Matrix-Free Two-to-Infinity and One-to-Two Norms Estimation

### v1

@@ 5.2 WideResNet Jacobian
To validate our methods on real-world data, we evaluate them on the Jacobian matrix $J\in\mathbb{R}^{3\cdot 32\cdot 32\times 100}$ of a WideResNet-16-10 (Zagoruyko and Komodakis 2016) trained on CIFAR-100 (Krizhevsky, Hinton et al. 2009). Given the structure of $J$ (the number of columns is much smaller than the number of rows), we expect that the $AA^{\top}P$ term (see Equation 1) in TwINEst++ will capture almost the entire span of $AA^{\top}$ with high probability. Therefore, TwINEst++ should outperform the basic TwINEst method on this matrix.
Figure 2 confirms our hypothesis: the TwINEst++ algorithm achieves rapid convergence consistent with the rank of matrix $J$, significantly outperforming other algorithms. Again, the Adaptive Power Method from (Roth, Kilcher, and Hofmann 2020) fails to converge. For a comprehensive ablation, we provide a comparison between the relative error of the methods and their floating point operation counts (FLOPs) in Figure 5.
@@ 5.3 Deep Learning Application
We study whether penalizing the two-to-infinity norm of the input-output Jacobian can improve the generalization ability and adversarial robustness of neural networks in image classification. Theorem 1 in (Roth, Kilcher, and Hofmann 2020) establishes a formal connection between $\ell_{p\to q}$ Jacobian norm regularization and adversarial training. This motivates our hypothesis that two‑to‑infinity regularization can enhance adversarial robustness of deep neural networks.
We compare our regularizer to the standard Jacobian-based penalties: Frobenius norm (Hoffman, Roberts, and Yaida 2019), spectral norm (Roth, Kilcher, and Hofmann 2020), and infinity norm ($\ell_{\infty}$) (Roth, Kilcher, and Hofmann 2020). For most image-classification datasets (e.g., CIFAR-100 (Krizhevsky, Hinton et al. 2009), TinyImageNet (Le and Yang 2015)), the number of output classes is much smaller than the number of input features (pixels). Consequently, the Jacobian of an image classifier trained on these datasets is a tall matrix. Hence, the two-to-infinity norm’s finer control over the Jacobian’s elements leads us to expect our regularizer to outperform standard norm penalties such as the Frobenius and spectral norms.
@@
Experiments are conducted on CIFAR-100 (Krizhevsky, Hinton et al. 2009) and TinyImageNet (Le and Yang 2015) datasets using the WideResNet-16-10 (Zagoruyko and Komodakis 2016) architecture, implemented in PyTorch. The hyperparameters are provided in Section D.4. We evaluate each method by reporting the final test accuracy, the stable rank of the Jacobian (computed as $\|J\|_{F}^{2}/\|J\|_{2}^{2}$), and adversarial metrics: accuracy after FGSM (Goodfellow, Shlens, and Szegedy 2015) and PGD (Madry et al. 2018) attacks with $2$ steps.
@@ 5.4 RecSys Application
Note that computing the full score matrix $\hat{R}$ at every training iteration is computationally prohibitive due to its large size and memory requirements. To address this, we approximate the two-to-infinity norm using our TwINEst algorithm. We compare our two-to-infinity regularizer with the standard weight decay used in the UltraGCN model, as well as with the two-to-infinity norm of the factors ($\|U\|_{2\to\infty}^{2}+\|V\|_{2\to\infty}^{2}$), which is closely related to the max-norm regularizer popular in recommender systems (Srebro, Rennie, and Jaakkola 2004).
Experiments are conducted on the MovieLens-1M (Harper and Konstan 2015), Yelp-2018 (Asghar 2016), and CiteULike (Wang, Chen, and Li 2013) datasets, using the UltraGCN architecture (Mao et al. 2021) implemented in PyTorch. The hyperparameter configurations and the dataset preprocessing pipeline are provided in Section D.7. We evaluate methods by reporting the NDCG@10 metric (which rewards ranking relevant items near the top) (Järvelin and Kekäläinen 2002). Additional results using other metrics are provided in Section D.8.
@@ 6 Conclusion
Acknowledgement
This research was supported in part through computational resources of HPC facilities at HSE University (Kostenetskiy, Chulkevich, and Kozyrev 2021).
@@ Appendix C Adaptive Power Method
Algorithm 3 Adaptive Power Method for Two-to-Infinity Norm from (Higham 1992; Roth, Kilcher, and Hofmann 2020)
@@ D.2 Singular Values of Evaluation Matrices
D.3 Additional Figures for Method Comparison
@@ D.5 Regularizers for Image Classification
For Frobenius-norm regularization, we use the algorithm introduced in (Hoffman, Roberts, and Yaida 2019), relying on $1$ Jacobian–vector multiplication. For spectral and infinity norm regularization, we adopt the algorithm described in (Roth, Kilcher, and Hofmann 2020), based on $3$ sequential Jacobian–vector multiplications. For two-to-infinity norm regularization, we employ the TwINEst algorithm, based on $5$ Jacobian–vector multiplications, $4$ of them executed in parallel.

### v2

@@ 5.2 WideResNet Jacobian
To validate our methods on real-world data, we evaluate them on the Jacobian matrix $J\in\mathbb{R}^{3\cdot 32\cdot 32\times 100}$ of a WideResNet-16-10 (Zagoruyko and Komodakis 2016) trained on CIFAR-100 (Krizhevsky et al. 2009). Given the structure of $J$ (the number of columns is much smaller than the number of rows), we expect that the $AA^{\top}P$ term (see Equation 1) in TwINEst++ will capture almost the entire span of $AA^{\top}$ with high probability. Therefore, TwINEst++ should outperform the basic TwINEst method on this matrix.
Figure 2 confirms our hypothesis: the TwINEst++ algorithm achieves rapid convergence consistent with the rank of matrix $J$, significantly outperforming other algorithms. Again, the Adaptive Power Method fails to converge. For a comprehensive ablation, we provide a comparison between the relative error of the methods and their floating point operation counts (FLOPs) in Figure 6.
@@ 5.3 Deep Learning Application
We study whether penalizing the two-to-infinity norm of the input-output Jacobian can improve the generalization ability and adversarial robustness of neural networks in image classification. Theorem 1 in (Roth et al. 2020) establishes a formal connection between $\ell_{p\to q}$ Jacobian norm regularization and adversarial training. This motivates our hypothesis that two‑to‑infinity regularization can enhance adversarial robustness of deep neural networks.
We compare our regularizer to the standard Jacobian-based penalties: Frobenius norm (Hoffman et al. 2019), spectral norm (Roth et al. 2020), and infinity norm ($\ell_{\infty}$) (Roth et al. 2020). For most image-classification datasets (e.g., CIFAR-100, TinyImageNet (Le and Yang 2015)), the number of output classes is much smaller than the number of input features (pixels). Consequently, the Jacobian of an image classifier trained on these datasets is a tall matrix. Hence, the two-to-infinity norm’s finer control over the Jacobian’s elements leads us to expect our regularizer to outperform standard norm penalties such as the Frobenius and spectral norms.
@@
Experiments are conducted on CIFAR-100 and TinyImageNet datasets using the WideResNet-16-10 architecture, implemented in PyTorch. The hyperparameters are provided in Section D.4. We evaluate each method by reporting the final test accuracy, the stable rank of the Jacobian (computed as $\|J\|_{F}^{2}/\|J\|_{2}^{2}$), and adversarial metrics: accuracy after FGSM (Goodfellow et al. 2015) and PGD (Madry et al. 2018) attacks with $2$ steps.
@@ 5.4 RecSys Application
Note that computing the full score matrix $\hat{R}$ at every training iteration is computationally prohibitive due to its large size and memory requirements. To address this, we approximate the two-to-infinity norm using our TwINEst algorithm. We compare our two-to-infinity regularizer with the standard weight decay used in the UltraGCN model, as well as with the two-to-infinity norm of the factors ($\|U\|_{2\to\infty}^{2}+\|V\|_{2\to\infty}^{2}$), which is closely related to the max-norm regularizer popular in recommender systems (Srebro et al. 2004).
Experiments are conducted on the MovieLens-1M (Harper and Konstan 2015), Yelp-2018 (Asghar 2016), and CiteULike (Wang et al. 2013) datasets, using the UltraGCN architecture implemented in PyTorch. The hyperparameter configurations and the dataset preprocessing pipeline are provided in Section D.7. We evaluate methods by reporting the NDCG@10 metric (which rewards ranking relevant items near the top) (Järvelin and Kekäläinen 2002). Additional results using other metrics are provided in Section D.8.
@@ 6 Conclusion
Acknowledgements
The work was supported by the grant for research centers in the field of AI provided by the Ministry of Economic Development of the Russian Federation in accordance with the agreement  [^00] C313925P4E0002 and the agreement with HSE University № 139-15-2025-009. This research was supported in part through computational resources of HPC facilities at HSE University (Kostenetskiy et al. 2021).
@@ Appendix C Adaptive Power Method
Algorithm 3 Adaptive Power Method for Two-to-Infinity Norm from (Higham 1992; Roth et al. 2020)
@@ D.2 Singular Values of Evaluation Matrices
D.3 Ablation for Method Comparison
Figure 5 presents an ablation study of different strategies for choosing $r$ (see Section 4). The base TwINEst++ algorithm uses a partition with $p=1/3$ ($r=p\cdot m=m/3$, where $m$ is the whole sampling budget). When $p\to 0$, TwINEst++ behaves like TwINEst, as expected, since the low-rank term has a negligible effect. When the low-rank term can capture the dominant eigenspace, as in Figure 5(c), it is better to use a greater $p$, as fewer terms will need to be approximated via the stochastic algorithm. Thus, $p$ represents a trade-off between not deviating too far from TwINEst in unfavorable settings and preserving most of the dominant eigenspace in the low-rank case.
(a) Synthetic data with $\Delta=10^{-2}$.
(b) Synthetic data with $\Delta=10^{-1}$.
(c) Jacobian matrix of WideResNet-16-10.
Figure 5: Comparison of different strategies for choosing $r$ in TwINEst++ algorithm. The plot shows the relative error versus number of matrix-vector multiplications, averaged over 500 trials.
@@ D.5 Regularizers for Image Classification
For Frobenius-norm regularization, we use the algorithm introduced in (Hoffman et al. 2019), relying on $1$ Jacobian–vector multiplication. For spectral and infinity norm regularization, we adopt the algorithm described in (Roth et al. 2020), based on $3$ sequential Jacobian–vector multiplications. For two-to-infinity norm regularization, we employ the TwINEst algorithm, based on $5$ Jacobian–vector multiplications, $4$ of them executed in parallel.